• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pzhu

    Log entry time 23:52:54 on March 04, 2012

    Entry number 363750

    keyword=swing shift summary

    18:30 finished harp check and VDC check, hall status from control access
    to beam permit
    19:40 harp is still not working,1H04 can see a very noise signal after
    exchanged the preamplifier,1H05A can not see any signal.we plan to do
    compton first
    11:50 some one changed the left arm's coda configuration which caused
    there is no bpm information from 2754(last optics run) to 2768. We will
    take the optics run later after compton run